OpenStudy (anonymous):

solve for ask X... ln(x+2)=10-ln(x+1) can anyone actually help me?

OpenStudy (anonymous):

hint: \[e ^{x+2}+e ^{x+1}=10\]

OpenStudy (anonymous):

\[e ^{x+2+x+1}= In10\]

OpenStudy (anonymous):

not unerstanding I went in the direction of ... ln(x+2)+ln(x+1)=10 \[\log_{e}\frac{ x+2 }{ x+1 }=10 \]